# Diff-Reg v1: Diffusion Matching Model for Registration Problem

### v3

@@ Abstract
Establishing reliable correspondences is essential for registration tasks such as 3D and 2D3D registration. Existing methods commonly leverage geometric or semantic point features to generate potential correspondences. However, these features may face challenges such as large deformation, scale inconsistency, and ambiguous matching problems (e.g., symmetry). Additionally, many previous methods, which rely on single-pass prediction, may struggle with local minima in complex scenarios. To mitigate these challenges, we introduce a diffusion matching model for robust correspondence construction. Our approach treats correspondence estimation as a denoising diffusion process within the doubly stochastic matrix space, which gradually denoises (refines) a doubly stochastic matching matrix to the ground-truth one for high-quality correspondence estimation. It involves a forward diffusion process that gradually introduces Gaussian noise into the ground truth matching matrix and a reverse denoising process that iteratively refines the noisy matching matrix. In particular, the feature extraction from the backbone occurs only once during the inference phase. Our lightweight denoising module utilizes the same feature at each reverse sampling step. Evaluation of our method on both 3D and 2D3D registration tasks confirms its effectiveness. The code is available at https://github.com/wuqianliang/Diff-Reg.
Index Terms:
3D Registration, 2D3D Registration, Diffusion model, Doubly Stochastic Matrix Space.
I Introduction
@@
The 3D registration problem, encompassing point cloud registration and image-to-point-cloud registration is critical in various computer vision and computer graphics applications, including 3D reconstructions, localization, and robotics. These applications usually necessitate precise correspondences (matchings) between point cloud pairs or image-to-point-cloud pairs for reliable rigid transformation or non-rigid deformation estimations.
@@
The goal of achieving accurate matchings is to identify the most significant correspondences[1, 2, 3] with local or global semantic or geometric consistency[4, 5, 6, 7]. However, this objective would be challenging especially in situations with globally ambiguous matching patches, large deformation, scale inconsistency, and low overlapping problems (refer to Fig.1).
@@
Recently, deep learning-based feature matching methods[6, 8, 9, 10, 11, 12, 2, 13, 4] have achieved significant progress in point cloud registration by employing UNet-like[14, 15] backbones to extract superpoints (subsampled patches) and their associated features. These methods typically compute an initial matching matrix between superpoints in the feature space. Additionally, outlier rejection techniques[16, 17, 18, 19, 20] propose specialized methodologies to identify improved inlier correspondences based on certain semantic or geometric priors[1, 7, 21, 22]. However, these methods usually rely on a single-pass prediction of correspondences, which may not always yield optimal results.
@@
In this paper, drawing inspiration from the diffusion model [23, 24, 25, 26], we introduce a diffusion matching model in the doubly stochastic matrix space[27] to progressively denoise (refine) the matching matrix to the optimal one for robust correspondence estimation. By training a diffusion model with the doubly stochastic matrix space as a feasible solution domain, we effectively learn a generalized optimization algorithm specifically designed for the doubly stochastic matrix space, adapted to the characteristics of the dataset or scene. Since the doubly stochastic matrix space is a convex set, the diffusion model can find the optimal solution by leveraging the learned ”direction”. Our diffusion matching model consists of two main components: a forward diffusion process and a reverse denoising process, which operates within the matrix space. The forward diffusion process gradually introduces Gaussian noise into the ground truth matching matrix, while the reverse denoising process iteratively refines the noisy matrix to the optimal one. For efficiency, we propose a novel and generalized lightweight denoising module that can be adapted to 2D2D image matching, 2D3D, and 3D registration tasks.
Finally, we establish a specific variational lower bound associated with our diffusion matching model in the doubly stochastic matrix space and a simplified version of the objective function to train our framework effectively.
@@
Why diffused in the doubly stochastic matrix space? Tasks such as image or point cloud registration face challenges like scale inconsistency, large deformation, ambiguous matching, and low overlapping. Several state-of-the-art (SOTA) studies[6, 13, 7, 28, 29, 17, 20] have attempted to encode high-order combinational geometric consistency. However, these manually crafted designs may not encompass all potential effective strategies for various challenging scenarios (e.g., large deformation). The diffusion process in the matrix space serves as a practical data augmentation technique that can generate additional training samples incorporating any-order combinational consistency, offering a promising approach to address these challenges.
@@
Our approach presents several advantages compared to previous registration methods. The forward diffusion process generates diverse training samples, acting as data augmentation for the feature backbone and single-pass prediction head, thus improving the model’s generalization capability. In contrast, the drawback of one-pass prediction methods is that if correspondences are predicted in a local minimum, subsequent outlier rejection or post-processing steps may face significant challenges. Our reverse denoising process, guided by the posterior distribution, allows for escaping from local minima, enabling the process to initiate from either white noise or any initial solution. Another enhancement is eliminating the feature backbone during the reverse denoising process at inference time. This streamlined design enables the denoising sampling process to explore a broader solution space (e.g., the matrix space), increasing diversity and facilitating more iterative steps. The findings from our empirical experiments support these claims.
@@
  • We conducted comprehensive experiments on the real-world 4DMatch[30], 3DMatch[31], and RGB-D Scenes V2[31, 4] datasets to validate the effectiveness of our diffusion matching model on 3D registration and 2D3D registration task.
(a) Large Deformation
(b) Symmetry
(c) Low Overlapping
(d) Large Motion and Low Overlapping
(e) Scale Inconsistency
Figure 1: Some challenges in 3D (and 2D3D) registration problem. (d) Presence of both low overlapping and large deformation issues simultaneously. (e) Losing the correspondence far from the camera. Please zoom in for details.
II Related work
II-A 3D and 2D3D Registration
The registration problem estimates the transformation between the point cloud pair or the image-to-point-cloud pair. Recently, there have been significant advancements in feature learning-based methods for point cloud registration. Many of these state-of-the-art approaches, such as [3, 11, 12, 6, 32, 33], leverage a backbone architecture similar to KPConv[14] to downsample points and generate features with larger receptive fields. To further enhance the performance of these methods, they integrate prior knowledge and incorporate learnable outlier rejection modules. For instance, GeoTR [6] introduces angle-wise and edge-wise embeddings into the transformer encoder, while RoITr [2] integrates local Point Pair Features (PPF) [34] to improve rotation invariance.
@@ II Related work
Recently, significant advancements have been made in 2D3D registration methods[36, 37, 38, 4]. These methods face similar challenges to 3D registration tasks, with the additional complexity of scale inconsistency caused by the perspective projection of images. To address the issue of scale inconsistency, we propose the incorporation of a pre-trained feature backbone, DINO v2 [39], which offers superior multiscale features. Additionally, implementing diffused data augmentation in our diffusion matching model can enhance the ability to identify prominent combinational and consistent correspondences.
II-B Diffusion Models for 3D Registration
Recently, the diffusion model[23, 40, 24] has made great development in many fields, including human pose estimation [41, 42], camera pose estimation [43], object detection[44], segmantation [45, 46]. These developments have been achieved through a generative Markov Chain process based on the Langevin MCMC[47] or a reversed diffusion process[24]. Recognizing the power of the diffusion model to iteratively approximate target data distributions from white noise using hierarchical variational decoders, researchers have started applying it to point cloud registration and 6D pose estimation problems.
@@ II-B Diffusion Models for 3D Registration
III The proposed Approach
III-A Problem Formulation
@@
Given source point clouds $P\in\mathbb{R}^{N{\times}3}$ and target point clouds $Q\in\mathbb{R}^{M{\times}3}$, the 3D registration task is to find top-k correspondences $\mathcal{C}$ from matching matrix $E$ and to conduct warping transformation ($T\in SE(3)$ for rigid transformations, and 3D flow fields for non-rigid transformations) to align the overlap region of $P$ and $Q$. In the context of 2D-3D registration, with a source image $X\in\mathbb{R}^{H{\times}W{\times}2}$ and target point cloud $Q\in\mathbb{R}^{M{\times}3}$, the standard pipeline involves determining the top-k correspondences $\mathcal{C}=\{(x_{i},q_{j})|x_{i}\in\mathbb{R}^{2},q_{j}\in\mathbb{R}^{3}\}$, and then estimating the rigid transformation $T\in SE(3)$ by minimizing the 2D projection error:
$\underset{T\in SE(3)}{\min}\sum\limits_{x_{i},q_{j}\in\mathcal{C}}||Proj(T(q_{j}),K)-x_{i}||_{2}$
@@
where $K$ represents the camera intrinsic matrix, and $Proj(\cdot,\cdot)$ denotes the projection function from 3D space to the image plane.
III-B Preliminaries
Revisiting Doubly Stochastic Matrix. We can represent the point clouds $P$ and $Q$ as two graphs, denoted as $\mathcal{G}_{1}=\left\{P,E^{P}\right\}$ and $\mathcal{G}_{2}=\left\{Q,E^{Q}\right\}$, where $E^{P}$ and $E^{Q}$ are respective edge sets. The matching matrix between these two graphs is a one-to-one mapping $E\in\{0,1\}^{N{\times}M}$. In cases where $N\neq M(e.g.,N>M)$, we can introduce $N-M$ dummy points in $Q$ to make a square matching matrix, also known as a permutation matrix =$\mathcal{M}=\left\{A:A1_{N}=1_{N},A^{T}1_{N}=1_{N},A\geq 0\right\}$. Then, we further employ sinkhorn iterations[50] to convert this non-negative real matrix into a “doubly stochastic” matrix, which has uniform row sum $M$ and column sum $N$ [27].
In the registration tasks, the many-to-one relationship between the matching matrix space and the warping operation space serves as a motivation for us:
$&\underset{E\in\mathcal{M}}{\min}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{M}E_{ij}*||W^{E}(p_{i})-q_{j}||_{2}\leq\\ &\mathop{\min}\limits_{W\in SE(3)/\mathbb{R}^{3N}/\mathbb{R}^{2N}}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{M}\hat{E}^{W}_{ij}*||W(p_{i})-q_{j}||_{2}\end{aligned}$
where $W^{E}$ is a warping generated by $E$ and $\hat{E}^{W}$ is a matching matrix computed by the warping W.
@@ III-B Preliminaries
Figure 2: Overview of our diffusion matching model. The forward diffusion process is driven by the Gaussian transition kernel $q(E^{t}|E^{t-1})$, which has a closed form $q(E^{t}|E^{0})$. The denoising model $g_{\theta}(E^{t})$ learns a reverse denoising gradient that points to the target solution $E^{0}$. During inference in the reverse sampling process, we utilize the predicted $\hat{E}_{0}$ and DDIM. ([24]) to sampling $E^{t-1}$.
III-C Overview
@@
Our framework comprises a feature backbone (e.g., KPConv[14] or ResNet[15]) and a diffusion matching model[23]. In the 3D registration task, the KPConv backbone takes source point clouds $P$ and target $Q$ as input and performs downsamplings to obtain the superpoints $\hat{P}$ and $\hat{Q}$, along with their associated features $F^{\hat{P}}\in\mathbb{R}^{N{\times}d}$ and $F^{\hat{Q}}\in\mathbb{R}^{M{\times}d}$. In the 2D3D registration task, a ResNet[15] with FPN[51] downsamples the image $X\in\mathbb{R}^{H{\times}W{\times}2}$ (in image-point-coud pair $(X,Q)$) to the superpixels $\hat{X}\in\mathbb{R}^{\hat{H}{\times}\hat{W}{\times}2}$ and give the associated image feature $F^{\hat{X}}\in\mathbb{R}^{\hat{H}{\times}\hat{W}{\times}d}$, while the target point cloud $Q$ is processed by a KPConv backbone, similar to the 3D registration task. Additionally, a depth map $D^{\hat{X}}$ and new superpixel feature $F^{\hat{X}}_{dino}$ for superpixels $\hat{X}$ are given by the pre-trained depth estimation model[52] and visual feature backbone DINO v2[39], respectively.
@@
Our diffusion module $g_{\theta}$ (refer to section III-E) takes these superpoints (or superpixels) and associated features as inputs. We employ one loss for our denoising module and another for the single-pass prediction head by leveraging the backbone features for supervision during joint training. During inference, we utilize $g_{\theta}$ in the reverse sampling process to predict the target matching matrix $E^{0}$ from any noisy one and apply DDIM [24] in the reverse sampling process to sample a more accurate matching matrix. The denoising module $g_{\theta}$ primarily consists of four key components: (1) Sinkhorn Projection[50], (2) Weighted SVD[53], (3) Denoising Transformer Network[54] and (4) Matching function.
III-D Diffusion Model in Doubly Stochastic Matrix space
@@
In this section, we introduce the construction of our diffusion matching model for generating the matching matrix between two scans. We denote the matching matrix as $E\in\{0,1\}^{N{\times}M}$, and we assume $E$ is defined in a nonsquare “doubly stochastic” matrix space $\mathcal{M}$.
@@
Forward Diffusion Process. As mentioned in DDPM[23], the forward diffusion process is fixed to a Markovian chain, denoted as $q(E^{1:T}|E^{0})$, which generates a sequence of latent variables $E^{t}$ by gradually injecting Gaussian noise into the ground truth matching matrix $E^{0}$. The diffused matching matrix $E^{t}$ at arbitrary timestep t has a closed form:
$E^{t}\sim q(E^{t}|E^{0})=N(E^{t};\sqrt{\bar{\alpha}}E^{0},(1-\bar{\alpha})\textbf{I})).$ (1)
@@
where the added noise over each element of the matrix is sampled independently and identically distributed (i.i.d.). However, this diffused $E^{t}\sim q(E^{t}|E^{0})$ is a continuous matrix in $R^{N\times M}$, which is outside the feasible solution space of matching matrices (i.e., doubly stochastic matrix manifolds). To address this issue, we apply the following projection to confine the matrix $E^{t}$ to the feasible solution space $\mathcal{M}$:
@@
Reverse Denoising Sampling Process. Given a diffusion Markovian chain $E^{0}\rightarrow E^{1}\rightarrow...\rightarrow E^{T}$, we need to learn a reverse transition kernel with the posterior distribution $q(E^{t-1}|E^{t},E^{0})$ to sample the reverse Markovian chain $E^{T}\rightarrow E^{T-1}\rightarrow...\rightarrow E^{0}$ from a white noise $E^{T}$ to achieve the target matching matrix. The posterior distribution $q(E^{t-1}|E^{t},E^{0})$ conditioned on $E^{0}$ and $E^{t}$ is defined as:
@@
$&q(E^{t-1}|E^{t},E^{0})=\frac{q(E^{t}|E^{t-1},E^{0})q(E^{t-1}|E^{0})}{q(E^{t}|E^{0})}\\ &\propto N(E^{t-1};\underbrace{\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})E^{t}+\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})E^{0}}{1-\bar{\alpha}_{t}}}_{\mu_{q}(E^{t},E^{0})},\\ &\underbrace{\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\textbf{I}}_{\Sigma_{q}(t)}).\end{aligned}$ (3)
For latent variable $E^{1:T}$, the Evidence Lower Bound (ELBO) for $E^{0}$ with distribution $q$ formulates as:
$&\log(p_{\theta}(E^{0}))\\ &\geq\mathbb{E}_{q(E^{1:T}|E^{0})}\left[\log\left(\frac{p_{\theta}(E^{0:T})}{q(E^{1:T}|E^{0})}\right)\right]=L_{vb}(E^{0})\\ &=\mathbb{E}_{q(E^{1}|E^{0})}\left[\log(p_{\theta}(E^{0}|E^{1}))\right]+\mathbb{E}_{q(E^{T}|E^{0})}\left[\log(\frac{p_{\theta}(E^{T})}{q(E^{T}|E^{0})})\right]\\ &+\sum_{t=2}^{T}\mathbb{E}_{q(E^{t},E^{t-1}|E^{0})}\left[\frac{p_{\theta}(E^{t-1}|E^{t})}{q(E^{t-1}|E^{t},E^{0})}\right]\\ &=C_{\theta}(E^{T},E^{1},E^{0})\\ &-\sum_{t=2}^{T}\underbrace{\mathbb{E}_{q(E^{t}|E^{0})}\left[D_{KL}\left(q(E^{t-1}|E^{t},E^{0})||p_{\theta}(E^{t-1}|E^{t})\right)\right]}_{\text{denoising matching term}}.\end{aligned}$
Since we propose a lightweight denoising network $g_{\theta}(E^{t})$ seek to predict $E^{0}$ from any noisy one $E^{t}$, the ELBO can be simplified to:
[…]
where $\sigma^{2}_{q}(t)=\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}}$.
Based on the derivation of Eqn. (4), we can utilize the following simplified version loss to train $p_{\theta}(E^{0}|E^{t})$:
@@
$&L_{simple}=\\ &-\mathbb{E}_{q(E^{0})}\left[\sum\limits_{t-1}^{T}\left(\frac{T-t+1}{T}\right)\mathbb{E}_{q(E^{t}|E^{0})}logp_{\theta}(E^{0}|E^{t})\right],\end{aligned}$ (5)
where the reweighting term $\frac{T-t+1}{T}$ implies that the loss at time steps closer to T is weighted less than at earlier steps.
III-E The Lightweight Denoising Module $g_{\theta}$

### v4

# Diff-Reg: Diffusion Model in Doubly Stochastic Matrix Space for Registration Problem

@@ Abstract
Establishing reliable correspondences is essential for 3D and 2D-3D registration tasks. Existing methods commonly leverage geometric or semantic point features to generate potential correspondences. However, these features may face challenges such as large deformation, scale inconsistency, and ambiguous matching problems (e.g., symmetry). Additionally, many previous methods, which rely on single-pass prediction, may struggle with local minima in complex scenarios. To mitigate these challenges, we introduce a diffusion matching model for robust correspondence construction. Our approach treats correspondence estimation as a denoising diffusion process within the doubly stochastic matrix space, which gradually denoises (refines) a doubly stochastic matching matrix to the ground-truth one for high-quality correspondence estimation. It involves a forward diffusion process that gradually introduces Gaussian noise into the ground truth matching matrix and a reverse denoising process that iteratively refines the noisy one. In particular, we deploy a lightweight denoising strategy during the inference phase. Specifically, once points/image features are extracted and fixed, we utilize them to conduct multiple-pass denoising predictions in the reverse sampling process. Evaluation of our method on both 3D and 2D-3D registration tasks confirms its effectiveness. The code is available at https://github.com/wuqianliang/Diff-Reg.
footnotetext: Corresponding Authors
1 Introduction
@@
The 3D registration problem, encompassing point cloud registration and image-to-point cloud registration, is critical in various computer vision and computer graphics applications, including 3D reconstructions, localization, AR, and robotics. These applications usually necessitate precise correspondences (matchings) between point cloud pairs or image-to-point cloud pairs for reliable rigid transformation or non-rigid deformation estimations.
@@
The goal of achieving accurate matchings is to identify the most significant correspondences [70, 66, 4] with local or global semantic or geometric consistency [25, 49, 40, 53]. However, this objective would be challenging, especially in situations with globally ambiguous matching patches, large deformation, scale inconsistency, and low overlapping problems.
@@
Recently, deep learning-based feature matching methods [40, 28, 63, 34, 19, 65, 66, 62, 25] have achieved significant progress in point cloud registration by employing UNet-like [45, 17] backbones to extract superpoints (subsampled patches) and their associated features. These methods typically compute an initial matching matrix between superpoints in the feature space. Additionally, outlier rejection techniques [60, 3, 10, 20, 69] propose specialized methodologies to identify improved inlier correspondences based on certain semantic or geometric priors [70, 53, 39, 14, 56, 59, 57, 58]. However, these methods usually rely on a single-pass prediction of correspondences, which may not always yield optimal results.
@@
In this paper, drawing inspiration from the diffusion model [18, 43, 48, 2], we introduce a diffusion matching model in the doubly stochastic matrix space[7]. By training a diffusion model with the doubly stochastic matrix space as a feasible solution domain, we effectively learn a generalized optimization algorithm specifically designed for the doubly stochastic matrix space, adapted to the characteristics of the dataset or scene. Our diffusion matching model consists of two main components: a forward diffusion process and a reverse denoising process, which operates within the matrix space. The forward diffusion process gradually introduces Gaussian noise into the ground truth matching matrix, while the reverse denoising process iteratively refines the noisy matrix to the optimal one. For efficiency, we propose a novel and generalized lightweight denoising module that can be adapted to 2D-3D and 3D registration tasks. Finally, we establish a specific variational lower bound associated with our diffusion matching model in the doubly stochastic matrix space and a simplified version of the objective function to train our framework effectively.
@@
Why diffused in the doubly stochastic matrix space? Tasks like image or point cloud registration face challenges like scale inconsistency, large deformation, ambiguous matching, and low overlapping. Several state-of-the-art studies [40, 62, 53, 54, 33, 3, 69] have attempted to encode high-order combinational geometric consistency. However, these manually crafted designs may not encompass all potential effective strategies for various challenging scenarios (e.g., large deformation). The diffusion process in the matrix space is a practical data augmentation technique that can generate additional training samples incorporating any-order combinational geometric consistency, offering a promising approach to address these challenges. A doubly stochastic matching matrix is, in fact, a dual-directional mapping. It offers a one-to-one mapping relation constraint for any kind of two-view (and any two-modality) matching/registration problem. The diffusion process within the matching matrix space naturally provides a broader range of training samples. Furthermore, the many-to-one relationship between the matching matrix space and the warping operation space can facilitate our method in learning a more optimal sampling path used in the reverse sampling process, in contrast to methods that conduct diffusion in SE(3) space.
@@
Our approach presents several advantages compared to previous registration methods. The forward diffusion process generates diverse training samples, acting as data augmentation for the feature backbone and single-pass prediction head. The drawback of single-pass prediction methods is that if correspondences are predicted in a local minimum, subsequent outlier rejection or post-processing steps may face significant challenges. In contrast, our reverse denoising process, guided by the posterior distribution, allows for escaping from local minima, enabling the process to initiate from either white noise or any initial solution. Another enhancement is eliminating the feature backbone during the reverse denoising process at inference time. This streamlined design enables the denoising sampling process to explore a broader solution space (e.g., the matrix space), increasing diversity and facilitating more iterative steps. The findings from our empirical experiments support these claims.
@@
  • We conducted comprehensive experiments on the real-world 4DMatch [29], 3DMatch [68], and RGB-D Scenes V2 [68, 25] datasets to validate the effectiveness of our diffusion matching model on 3D registration and 2D-3D registration task.
2 Related work
2.1 3D and 2D-3D Registration
The registration problem estimates the transformation between the point cloud or image-to-point-cloud pair. Recently, there have been significant advancements in feature learning-based methods for point cloud registration. Many of these state-of-the-art approaches, such as [4, 19, 65, 40, 64, 67], leverage a backbone architecture similar to KPConv [45] to downsample points and generate features with larger receptive fields. To further enhance the performance of these methods, they integrate prior knowledge and incorporate learnable outlier rejection modules. For instance, GeoTR [40] introduces angle-wise and edge-wise embeddings into the transformer encoder, while RoITr [66] integrates local Point Pair Features (PPF) [13] to improve rotation invariance.
@@ 2 Related work
Recently, significant advancements have been made in 2D-3D registration methods [24, 50, 51, 25]. These methods face similar challenges to 3D registration tasks, with the additional complexity of scale inconsistency caused by the perspective projection of images. To address the issue of scale inconsistency, we propose the incorporation of a pre-trained feature backbone, DINO v2 [36], which offers superior multiscale features. Additionally, implementing diffused data augmentation in our diffusion matching model can enhance the ability to identify prominent combinational and consistent correspondences.
2.2 Diffusion Models for 3D Registration
Recently, the diffusion model [18, 44, 43] has made great development in many fields, including human pose estimation [15, 42], camera pose estimation [52], object detection [8], segmentation [5, 16]. These developments have been achieved through a generative Markov Chain process based on the Langevin MCMC [37] or a reversed diffusion process [43]. Recognizing the power of the diffusion model to iteratively approximate target data distributions from white noise using hierarchical variational decoders, researchers have started applying it to point cloud registration and 6D pose estimation problems.
@@ 2.2 Diffusion Models for 3D Registration
3 The proposed Approach
3.1 Problem Formulation
@@
Given source point clouds $\mathbf{P}\in\mathbb{R}^{N{\times}3}$ and target point clouds $\mathbf{Q}\in\mathbb{R}^{M{\times}3}$, the 3D registration task is to find top-k correspondences $\mathcal{C}$ from matching matrix $\mathbf{E}$ and to conduct warping transformation ($\mathbf{\Gamma}\in$ SE(3) for rigid transformations, and 3D flow fields for non-rigid transformations) to align the overlap region of $\mathbf{P}$ and $\mathbf{Q}$. In the context of 2D-3D registration, with a source image $\mathbf{X}\in\mathbb{R}^{H{\times}W{\times}2}$ and target point cloud $\mathbf{Y}\in\mathbb{R}^{M{\times}3}$, the standard pipeline involves determining the top-k correspondences $\mathcal{C}=\{(\mathbf{x}_{i},\mathbf{y}_{j})|\mathbf{x}_{i}\in\mathbb{R}^{2},\mathbf{y}_{j}\in\mathbb{R}^{3}\}$, and then estimating the rigid transformation $\mathbf{\Gamma}\in$ SE(3) by minimizing the 2D projection error:
$\underset{\mathbf{\Gamma}\in SE(3)}{\min}\sum\limits_{\mathbf{x}_{i},\mathbf{y}_{j}\in\mathcal{C}}||Proj(\mathbf{\Gamma}(\mathbf{y}_{j}),\mathbf{K})-\mathbf{x}_{i}||_{2}$
@@
where $\mathbf{K}$ represents the camera intrinsic matrix, and $Proj(\cdot,\cdot)$ denotes the projection function from 3D space to the image plane.
@@
Figure 1: Overview of our diffusion matching model. The forward diffusion process is driven by the Gaussian transition kernel $q(\mathbf{E}^{t}|\mathbf{E}^{t-1})$, which has a closed form $q(\mathbf{E}^{t}|\mathbf{E}^{0})$. The denoising model $g_{\theta}(\mathbf{E}^{t})$ learns a reverse denoising gradient that points to the target solution $\mathbf{E}^{0}$. During inference, in the reverse sampling process, we utilize the predicted $\hat{\mathbf{E}}_{0}$ and DDIM [43] to sampling $\mathbf{E}^{t-1}$.
3.2 Overview
@@
Our framework comprises a feature backbone (e.g., KPConv [45]/ResNet [17]) and a diffusion matching model [18]. In the 3D registration task, the KPConv backbone takes source point clouds $\mathbf{P}$ and target $\mathbf{Q}$ as input and performs downsamplings to obtain the superpoints $\hat{\mathbf{P}}$ and $\hat{\mathbf{Q}}$, along with their associated features $\mathbf{F}^{\hat{\mathbf{P}}}\in\mathbb{R}^{N{\times}d}$ and $\mathbf{F}^{\hat{\mathbf{Q}}}\in\mathbb{R}^{M{\times}d}$. In the 2D-3D registration task, a ResNet [17] with FPN [32] downsamples the image $\mathbf{X}\in\mathbb{R}^{H{\times}W{\times}2}$ (in image-point-coud pair $(\mathbf{X},\mathbf{Q})$) to the superpixels $\hat{\mathbf{X}}\in\mathbb{R}^{\hat{H}{\times}\hat{W}{\times}2}$ and give the associated image feature $\mathbf{F}^{\hat{\mathbf{X}}}\in\mathbb{R}^{\hat{H}{\times}\hat{W}{\times}d}$, while the target point cloud $\mathbf{Q}$ is processed by a KPConv backbone, similar to the 3D registration task. Additionally, a depth map $\mathbf{D}^{\hat{\mathbf{X}}}$ and new superpixel feature $\mathbf{F}^{\hat{\mathbf{X}}}_{dino}$ for superpixels $\hat{\mathbf{X}}$ are given by the pre-trained depth estimation model [61] and visual feature backbone DINO v2 [36], respectively.
@@
Our diffusion module $g_{\theta}$ (refer to Section 3.4) takes these superpoints (or superpixels) and associated features as inputs. We employ one loss for our denoising module and another for the single-pass prediction head during the training stage. During inference, we utilize $g_{\theta}$ in the reverse sampling process to predict the target matching matrix $\mathbf{E}^{0}$ from any noisy one and exploit DDIM [43] to sample a more accurate matching matrix. The denoising module $g_{\theta}$ primarily consists of four components: (1) Sinkhorn Projection [12], (2) Weighted SVD [6], (3) Warping Function (4) Denoising Transformer Network [47] and (5) Matching function. More details of our framework can be found in section 3.4 and the appendix.
3.3 Diffusion Model in Doubly Stochastic Matrix Space
@@
In this section, we introduce the construction of our diffusion matching model for generating the matching matrix between two scans. We denote the matching matrix as $\mathbf{E}\in\{0,1\}^{N{\times}M}$, and we assume $\mathbf{E}$ is defined in a nonsquare “doubly stochastic” matrix space $\mathcal{M}$ (refer to Appendix).
@@
Forward Diffusion Process. As mentioned in DDPM [18], the forward diffusion process is fixed to a Markovian chain, denoted as $q(\mathbf{E}^{1:T}|\mathbf{E}^{0})$, which generates a sequence of latent variables $\mathbf{E}^{t}$ by gradually injecting Gaussian noise into the ground truth matching matrix $\mathbf{E}^{0}$. The diffused matching matrix $\mathbf{E}^{t}$ at arbitrary timestamp t has a closed form:
$\mathbf{E}^{t}\sim q(\mathbf{E}^{t}|\mathbf{E}^{0})=\mathcal{N}(\mathbf{E}^{t};\sqrt{\bar{\alpha}}\mathbf{E}^{0},(1-\bar{\alpha})\textbf{I})).$ (1)
@@
where the added noise over each element of the matrix is sampled independently and identically distributed (i.i.d.). However, this diffused $\mathbf{E}^{t}\sim q(\mathbf{E}^{t}|\mathbf{E}^{0})$ is a continuous matrix in $\mathbb{R}^{N\times M}$, which is outside the feasible solution space of matching matrices (i.e., doubly stochastic matrix manifolds). To address this issue, we apply the following projection to confine the matrix $\mathbf{E}^{t}$ to the feasible solution space $\mathcal{M}$:
@@
Reverse Denoising Sampling Process. Given a diffusion Markovian chain $\mathbf{E}^{0}\rightarrow\mathbf{E}^{1}\rightarrow...\rightarrow\mathbf{E}^{T}$, we need to learn a reverse transition kernel with the posterior distribution $q(\mathbf{E}^{t-1}|\mathbf{E}^{t},\mathbf{E}^{0})$ to sample the reverse Markovian chain $\mathbf{E}^{T}\rightarrow\mathbf{E}^{T-1}\rightarrow...\rightarrow\mathbf{E}^{0}$ from a white noise $\mathbf{E}^{T}$ to achieve the target matching matrix. The posterior distribution $q(\mathbf{E}^{t-1}|\mathbf{E}^{t},\mathbf{E}^{0})$ conditioned on $\mathbf{E}^{0}$ and $\mathbf{E}^{t}$ is defined as:
@@
$\begin{split}&q(\mathbf{E}^{t-1}|\mathbf{E}^{t},\mathbf{E}^{0})=\frac{q(\mathbf{E}^{t}|\mathbf{E}^{t-1},\mathbf{E}^{0})q(\mathbf{E}^{t-1}|\mathbf{E}^{0})}{q(\mathbf{E}^{t}|\mathbf{E}^{0})}\\ &\propto\mathcal{N}(\mathbf{E}^{t-1};\underbrace{\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\mathbf{E}^{t}+\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})\mathbf{E}^{0}}{1-\bar{\alpha}_{t}}}_{\mu_{q}(\mathbf{E}^{t},\mathbf{E}^{0})},\underbrace{\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\textbf{I}}_{\Sigma_{q}(t)}).\end{split}$ (3)
To effectively train our denoising network, we derive the variational lower bound of the log-likelihood of the training samples $\mathbf{E}^{0}$:
@@
$logp(\mathbf{E}^{0})\geq\mathbb{E}_{\mathbf{E}^{1:T}\sim q(\mathbf{E}^{1:T}|\mathbf{E}^{0})}\left[log\left(\frac{p_{\theta}(\mathbf{E}^{0:T})}{q(\mathbf{E}^{1:T}|\mathbf{E}^{0})}\right)\right]\propto\mathbb{E}_{q}\left[\sum\limits_{t=2}^{T}log(p_{\theta}(\mathbf{E}^{0}|\mathbf{E}^{t}))\right].\end{aligned}$ (4)
Based on the derivation of Eqn. (4), we can further simplify the variational lower bound above to train $p_{\theta}(\mathbf{E}^{0}|\mathbf{E}^{t})$:
$L_{simple}=-\mathbb{E}_{q(\mathbf{E}^{0})}\left[\sum\limits_{t-1}^{T}\mathbb{E}_{q(\mathbf{E}^{t}|\mathbf{E}^{0})}logp_{\theta}(\mathbf{E}^{0}|\mathbf{E}^{t})\right].$ (5)
3.4 The Lightweight Denoising Module $g_{\theta}$